# Zeroth-order Informed Fine-Tuning for Diffusion Model: A Recursive Likelihood Ratio Optimizer

### v2

@@ Abstract
The probabilistic diffusion model (DM), generating content by inferencing through a recursive chain structure, has emerged as a powerful framework for visual generation. After pre-training on enormous unlabeled data, the model needs to be properly aligned to meet requirements for downstream applications. How to efficiently align the foundation DM is a crucial task. Contemporary methods are either based on Reinforcement Learning (RL) or truncated Backpropagation (BP). However, RL and truncated BP suffer from low sample efficiency and biased gradient estimation respectively, resulting in limited improvement or, even worse, complete training failure. To overcome the challenges, we propose the Recursive Likelihood Ratio (RLR) optimizer, a zeroth-order informed fine-tuning paradigm for DM. The zeroth-order gradient estimator enables the computation graph rearrangement within the recursive diffusive chain, making the RLR’s gradient estimator an unbiased one with the lower variance than other methods. We provide theoretical guarantees for the performance of the RLR. Extensive experiments are conducted on image and video generation tasks to validate the superiority of the RLR. Furthermore, we propose a novel prompt technique that is natural for the RLR to achieve a synergistic effect. See our implementation at https://github.com/RTkenny/RLR-Opimtizer.
Machine Learning, ICML
Figure 1: Image and video samples generated from model fine-tuned by our RLR optimizer. More qualitative examples are included in Appendices A.3 and A.4.
@@ 1 Introduction
Probabilistic diffusion model (DM) (Sohl-Dickstein et al., 2015; Ho et al., 2020; Zhang et al., 2024b) has emerged as a transformative framework in high-fidelity data generation, demonstrating unmatched capabilities in diverse applications such as image synthesis (Podell et al., 2023), video generation (Wang et al., 2023a), and multi-modal data modeling. These models operate by iteratively denoising noisy data representations, effectively capturing complex underlying distributions. Despite their success, even fine-tuning DMs remains a daunting challenge due to the computational and caching demands associated with gradient computation during training, especially for large-scale tasks (Clark et al., 2023). This challenge has limited the broader deployment of DM in dynamic and resource-constrained environments.
@@
It is a natural way to fine-tune DMs via full backpropagation (BP) through all time steps (Rumelhart et al., 1986), which is theoretically functional, providing precise gradient estimation over the entire diffusion chain. However, the computational and memory overhead of BP scales prohibitively with the model size and the number of diffusion steps (Prabhudesai et al., 2023; Yuan et al., 2024; Clark et al., 2023; Prabhudesai et al., 2024), making full BP impractical for most real-world scenarios. Thus, truncating recursive differentiation becomes a common practice to alleviate memory constraints. But truncated BP suffers from structural bias, as it terminates gradient computation before sourcing to the input, only considering a limited subset of the diffusion chain. Fine-tuning based on a biased gradient estimation can inadvertently impair the optimization performance, resulting in model collapse, i.e., contents generated reducing to pure noise. Moreover, truncated BP fails to capture the multi-scale information across all time steps due to the absence of differentiation on early steps. The hierarchical feature of generation imposes a thorough gradient evaluation to retain fidelity from pixel to structural level.
@@
Reinforcement learning (RL) (Schulman et al., 2017) as a gradient computation trick has enabled another branch of DM fine-tuning (Lee et al., 2023; Black et al., 2023; Wallace et al., 2024; Fan et al., 2024). It typically ignores the differentiable connection between steps and recovers the gradient by estimation. RL avoids caching intermediate activations, significantly reducing memory requirements. It also supports gradient computation in a divided manner under extreme circumstances to accommodate insufficient memory. The cost to pay is the high variance of the estimated gradient. Even if the estimator is unbiased, the variance can result in wild sample-inefficient updates demonstrated by the slow convergence during training.
These limitations of BP and RL underscore the necessity of a more efficient, scalable, and stable fine-tuning approach that harmonizes computational tractability with optimization efficacy. To address this, we first investigate the recursive architecture of the DM and its relationship with perturbation-based gradient estimation. Subsequently, we conduct a rigorous analysis of the bias and variance properties of various estimation techniques, motivating our integrated optimizer.
Informed by perturbation-based analysis using Likelihood Ratio (LR) techniques, we propose the Recursive Likelihood Ratio (RLR) optimizer, a novel approach designed to address these challenges. The optimizer combines advantages from BP and RL but alleviates their shortcomings. By utilizing the inherent noise in the DM, a local computational chain is enabled, which can reduce the variance and better capture the multi-scale information. The gradient estimator from the RLR optimizer is unbiased and has low variance under the same computation budgets as other methods. Specifically, RLR manages memory overhead by limiting BP to strategically chosen segments of the diffusion chain while leveraging perturbation-based methods to estimate gradients for the remaining steps. This hybrid approach mitigates the structural bias of truncated BP and the high variance of RL, enabling stable and effective fine-tuning. Our contributions are threefold:
  • We give a thorough analysis of the bias and variance of different estimators in the recursive structure of DM, showing that the inherent noise in the DM can be used to perform computational graph rearrangement.
  • Based on the theoretical and empirical insight, we formulate the RLR optimizer which has desirable theoretical properties regarding bias and variance, and the convergence of the RLR optimizer is also guaranteed.
@@
  • Extensive evaluation of Text2Image and Text2Video tasks are conducted. Our RLR surpasses all the baselines by a large margin. Furthermore, we propose a prompt technique for our RLR, validating the applicability of our method.
2 Related works
Diffusion Probabilistic Models.
Denoising Diffusion Model (Ho et al., 2020) is one of the strongest models for generation tasks, especially for visual generation (Rombach et al., 2022; Peebles & Xie, 2023; Chen et al., 2024a). Extensive research has been conducted from theoretical and empirical perspectives (Song et al., 2020a; Karras et al., 2022; Song et al., 2020b). It has achieved phenomenal success in muti-modality generation, including image, video, audio, and 3D shapes. The DM is trained on enormous images and videos from the internet (Bain et al., 2021; Wang et al., 2023b; Schuhmann et al., 2022). Empowered by modern architecture (Vaswani, 2017), it has powerful learning capability for Pixel Space Distribution.
Alignment and Post-training.
After pre-training to learn the distribution of the targeted modality (Achiam et al., 2023; Kaplan et al., 2020), post-training is conducted to align the model toward specific preferences or tune the model to optimize a particular objective. RL has been utilized to align the foundation models toward various objectives (Ziegler et al., 2019; Lambert et al., 2022; Black et al., 2023). Supervised learning can also be applied to the post-training phase (Rafailov et al., 2024), either optimizing an equivalent objective (Wallace et al., 2024) or directly differentiating the reward model (Clark et al., 2023; Prabhudesai et al., 2023, 2024). For DM, most methods use a neural reward model to align the pre-trained model, and there has been a continual effort to design better reward models (He et al., 2024; Xu et al., 2024a, b).
Forward Learning Methods.
After extensive exploration of model training using forward inferences only (see, e.g., Peng et al., 2022; Hinton, 2022), forward-learning methods (Salimans et al., 2017; Malladi et al., 2023; Chen et al., 2023; Jiang et al., 2024) based on stochastic gradient estimation have recently emerged as a promising alternative to classical BP for large-scale machine learning problems (Zhao et al., 2024; Zhang et al., 2024a). Subsequent research (Ren et al., 2024; Chen et al., 2024b) has further optimized computational and storage overhead from various perspectives, achieving greater efficiency.
@@ 2 Related works
3 Diffusion Model under Perturbation-based Analysis
3.1 Recursive Structure in Diffusion Model.
DMs constituted of the forward process and the backward process work by injecting noises or perturbations into a recursive computation chain to generate data from a given distribution. By analyzing the final output and the inherent intermediate perturbation, a forward learning estimator is a free byproduct without “wasting the noise”.
In the forward process, data points (e.g., images) are gradually corrupted with Gaussian noise over a series of time steps, resulting in a sequence of noisy samples. The forward process is represented by a stochastic transformation $q(\bm{x}_{t}|\bm{x}_{0})$, where $\bm{x}_{0}$ is the original data, and $\bm{x}_{t}$ is the noisy sample at time $t$. The ultimate goal is for the model to learn the reverse process, starting from pure noise $\bm{x}_{T}$ and progressively recovering the original data $\bm{x}_{0}$.
The training objective is to learn the reverse diffusion process by predicting the noise added at each time step. This is achieved by minimizing the difference between the predicted noise and the actual noise added in the forward process. The objective is defined as follows:
$\min_{\theta}\mathbb{E}_{q(x_{0}),t\sim\mathcal{U}(0,T),q(\mathbf{x}_{t}|x_{0})}\left[\|\epsilon_{\theta}(x_{t},t)-\epsilon\|^{2}\right],$ (1)
where $\epsilon_{\theta}(x_{t},t)$ is the predicted noise at time $t$, and $\epsilon$ is the actual noise added during the forward process. The model’s neural network $\epsilon_{\theta}$ is trained to accurately predict this noise, enabling it to progressively denoise data during generation.
The backward process of the DM can be generally expressed by a stochastic differential equation (Karras et al., 2022; Song et al., 2020b):
$d\bm{x}_{t}=(h(\bm{x}_{t},t)+g(t)\epsilon_{\theta}(\bm{x}_{t},t))dt+d\bm{w}_{t},$ (2)
where $h(\bm{x}_{t},t)$ and $g(t)$ have analytical forms, and $d\bm{w}_{t}$ is the stochastic drift term. So the discretization to Equation (2) can be formulated as
$\bm{x}_{t-1}=\varphi_{t}(\bm{x}_{t};\theta)+\bm{z}_{t},$ (3)
where $\varphi_{t}(\bm{x}_{t};\theta)$ denotes the deterministic part of the discretization, and $\bm{z}_{t}=\sigma_{t}\epsilon$ stands for the stochastic part for $1\leq t\leq T$. Consider a general form of the sampling procedure:
$\bm{x}_{t-1}=\phi_{t}(\bm{x}_{t},\bm{z}_{t};\theta)=\varphi_{t}(\bm{x}_{t};\theta)+\sigma_{t}\epsilon.$ (4)
Then, when the backward process is deterministic, we define the diffusion chain as
$\bm{x}_{0}=\phi_{1:T}(\bm{x}_{T},\bm{z}_{1:T};\theta)=\phi_{1}\circ...\circ\phi_{T}(\bm{x}_{T},\bm{z}_{1:T};\theta)=$ (5)
$\phi_{1}(\phi_{2:T}(\bm{x}_{T},\bm{z}_{2:T};\theta),\bm{z}_{1};\theta)=\phi_{1:T-1}(\phi_{T}(\bm{x}_{T},\bm{z}_{T};\theta),$
$\bm{z}_{1:T-1};\theta)=\phi_{1}(\phi_{2:T-1}(\phi_{T}(\bm{x}_{T},\bm{z}_{T};\theta),\bm{z}_{2:T-1};\theta),\bm{z}_{1};\theta).$
After the pre-training phase, we need to conduct post-training to align the model to generate visually pleasing content. A reward model $R(\cdot)$ is deployed to evaluate the quality of the generated content $\bm{x_{0}}$. Therefore, we aim to solve the optimization problem for the post-training:
$\max_{\theta}\mathbb{E}[R(\bm{x}_{0})]=\max_{\theta}\mathbb{E}_{\bm{z}_{1:T}}[R(\phi_{1:T}(\bm{x}_{T},\bm{z}_{1:T};\theta))].$ (6)
Typically, a gradient estimator of the expected reward function is required for the optimization problem.
3.2 Perturbation-based Gradient Estimators
The LR method, also called the score function method, is a perturbation-based gradient estimation approach used in neural network training (Jiang et al., 2024; Ren et al., 2024). It provides an alternative to BP by bypassing recursive computations for the backward chain, making it computationally friendly and memory efficient. Especially, when applied to differentiate a very deep network or long time horizon of RNN, LR serves as an efficient surrogate gradient estimator for BP. There are two types of LR estimators, both of which have been proven to be unbiased in the literature (Glasserman, 1990; Jiang et al., 2024).
Half-Order (HO) Estimator
The first version borrows the intermediate noise as the perturbation to perform gradient estimation in Equation (3): $\bm{x}_{t-1}=\phi_{t}(\bm{x}_{t},\bm{z}_{t};\theta)=\varphi(\bm{x}_{t};\theta)+\bm{z}_{t}$, where $\bm{z}_{t}$ is the inherent noise of the DM following Gaussian distribution: $\bm{z}_{t}=\sigma_{t}\epsilon\sim\mathcal{N}(0,\sigma_{t}\bm{I})$. The gradient estimator for the first version takes the form:
$R(\bm{x}_{0})\sum_{j}D_{\theta}^{\top}\phi_{jh:jh+h-1}(\bm{x}_{jh},\bm{z}_{jh:jh+h-1};\theta)\nabla\ln f(\bm{z}_{jh}),$ (7)
where $f(\cdot)$ denotes the noise density, and $D_{\theta}\phi\in\mathbb{R}^{d_{\phi}\times d_{\theta}}$ denotes the Jacobian matrix of $\phi\in\mathbb{R}^{d_{\phi}}$ with respect to $\theta\in\mathbb{R}^{d_{\theta}}$. Since this estimator enables a local differentiation chain, $D_{\theta}^{\top}\phi_{jh:jh+h-1}(\bm{x}_{jh},\bm{z}_{jh:jh+h-1};\theta)$, of length $h$ after the perturbation $\bm{z}_{jh}$, it is a combination of Zeroth- and First-Order optimization, which we refer to as a Half-Order estimator.
Moreover, RL in DM training is fundamentally an HO method with a local chain length of 1 (Jiang et al., 2024). Action sampling can also be interpreted as perturbations applied to the model’s latent variables. However, it encounters challenges such as low sample efficiency and high variance in gradient estimation, as it disregards on-hand information along the forward computation.
Zeroth-Order (ZO) Estimator
The second version of LR directly injects exogenous noise to the parameters of the targeted module, $\bm{x}_{t-1}=\phi_{t}(\bm{x}_{t},\bm{z}_{t};\theta)=\varphi(\bm{x}_{t};\theta+\bm{z}_{t}).$ The estimator is then given by
$\sum_{t=1}^{T}R(\bm{x}_{0})\nabla\ln f(\bm{z}_{t}),$ (8)
which is commonly identified as the Zeroth-Order estimator. It is completely blind to all structural information from the forward computation, relying solely on output evaluations to estimate gradients. The well-known Simultaneous Perturbation Stochastic Approximation (SPSA) is covered by this version, when the antithetic variable trick is applied to perform randomized finite differences (Ren et al., 2024).
First-Order (FO) Estimator
As a contrasting extreme, BP is an FO method that requires the entire computational process to be pathwise differentiable. It leverages the chain rule of differentiation to compute gradients through the diffusive chain and the reward model, expressed as
$\nabla_{\theta}R(\bm{x}_{0})=$ $\bigg{[}\frac{\partial\phi_{1}(\bm{x}_{1},\bm{z}_{1};\theta)}{\partial\theta}+\sum_{i=2}^{T}\frac{\partial\phi_{i}(\bm{x}_{i},\bm{z}_{i};\theta)}{\partial\theta}$ (9)
$\prod_{j=1}^{i-1}\frac{\partial\phi(\bm{x}_{j},\bm{z}_{j};\theta)}{\partial\bm{x}_{j}}\bigg{]}^{\top}\frac{dR(\bm{x}_{0})}{d\bm{x_{0}}}.$
This approach ensures efficient gradient computation by utilizing the differentiable structure of the model.
The unbiasedness of the above estimators is provided in subsequent Theorem 5.1. We will compare their properties and then build a new RLR estimator based on their different merits in the next section.
4 Bias and Variance of Gradient Estimators
From the perspective of stochastic gradient estimation, there are three types of estimators for fine-tuning DM. BP and RL are covered by the aforementioned estimators, utilizing the differentiability or the inherent noise in the DM. We will further analyze the bias and variance of gradient estimations, motivating the formulation of the RLR optimizer.
4.1 Bias Analysis
The stochastic nature of the DM results in the variance of the FO estimator. As expected, the variance of the FO estimator is lower than that of the HO and ZO estimators because the differentiation leverages the precise structure of the neural network.
Performing full BP over the diffusive chain during training is computationally infeasible due to the large memory requirements. For example, training Stable Diffusion 1.4 with a batch size of 1 and 50 time steps would require approximately 1TB of GPU RAM (Prabhudesai et al., 2023).
To alleviate the memory burden of full-step BP, the truncated variant is often employed, where the total number of time steps $T$ in Equation (9) is replaced with a smaller surrogate $T^{\prime}\ll T$. However, the truncation introduces a structural bias into the gradient estimator. We have the following proposition to justify this structural bias.
Proposition 4.1.
Assume $R$ and $\phi$ are differentiable almost everywhere, $R$ and $\phi$ have bounded gradients or Lipschitz constants, then the FO estimator is unbiased. However, the truncate BP estimator $\nabla_{\theta}R(\bm{x}_{0})_{\mathrm{truncated}}$ has a structural bias, which can be specified as below:
$\nabla_{\theta}\mathbb{E}[R(\bm{x}_{0})]-\mathbb{E}[\nabla_{\theta}R(\bm{x}_{0})_{\mathrm{truncated}}]=\mathbb{E}_{z_{1:T}}$
$\bigg{[}\bigg{(}\sum_{i=T^{\prime}-1}^{T}\frac{\partial\phi(\bm{x}_{i};\theta)}{\partial\theta}\prod_{j=1}^{i-1}\frac{\partial\phi(\bm{x}_{j};\theta)}{\partial\bm{x}_{j}}\bigg{)}^{\top}\frac{dR(\bm{x}_{0})}{d\bm{x_{0}}}\bigg{]},$
where $\phi(\bm{x}_{j};\theta)$ denotes $\phi_{j}(\bm{x}_{j},z_{j};\theta)$ for simplicity.
The bias in the estimator can lead to numerical instability, resulting in suboptimal updates or even training failure, as the truncated gradient may not be a descent direction. The undesired consequences are as follows.
  • Model collapse: The model may fail to converge, degenerating into generating pure noise due to unstable optimization.
  • Failure to capture multi-scale information: By limiting the gradient computation to fewer time steps, truncated BP loses essential multi-scale visual information along the diffusive chain.
We conduct empirical studies to investigate the effects of truncation. As shown in Figure 3, truncated gradients result in a significant drop in reward scores during training; the fewer the truncated time steps, the more severe the model collapse. In our subsequent Diffusive Chain-of-Thought experiment, we demonstrate the enhanced capability of the proposed RLR optimizer to capture multi-scale information for visual generation, effectively mitigating the limitations of the truncated BP variant.
4.2 Variance Analysis
We know that FO, HO and ZO estimators are all unbiased. However, BP introduces significant computational and storage overhead. The following proposition demonstrates that this additional cost is, to some extent, justified, as BP leverages accurate internal structural to reduce estimation variance, and therefore the variance of BP is lower than zeroth-order methods.
Proposition 4.2.
Under Assumptions (A.1-3) in the Appendix, the variance of FO estimators in Equation (9) is less or equal to ZO estimators in Equation (8), i.e.
$\mathrm{Var}(\nabla_{\theta}R(\bm{x}_{0}))\leq\mathrm{Var}(R(\bm{x}_{0})\nabla\ln f(\bm{z})).$ (10)
Additionally, it is straightforward to conclude that the variance of HO estimator is also less or equal to that of ZO estimator, as it is essentially an FO estimator with a perturbation at the end of the chain. This argument is also discussed in (Salimans et al., 2017). The following Table 1 presents all the gradient computation methods discussed above.
Table 1: Comparison of different gradient estimators
Methods Unbiasedness Variance Memory

### v3

# Half-order Fine-Tuning for Diffusion Model: A Recursive Likelihood Ratio Optimizer

@@ Abstract
The probabilistic diffusion model (DM), generating content by inferencing through a recursive chain structure, has emerged as a powerful framework for visual generation. After pre-training on enormous data, the model needs to be properly aligned to meet requirements for downstream applications. How to efficiently align the foundation DM is a crucial task. Contemporary methods are either based on Reinforcement Learning (RL) or truncated Backpropagation (BP). However, RL and truncated BP suffer from low sample efficiency and biased gradient estimation, respectively, resulting in limited improvement or, even worse, complete training failure. To overcome the challenges, we propose the Recursive Likelihood Ratio (RLR) optimizer, a Half-Order (HO) fine-tuning paradigm for DM. The HO gradient estimator enables the computation graph rearrangement within the recursive diffusive chain, making the RLR’s gradient estimator an unbiased one with lower variance than other methods. We theoretically investigate the bias, variance, and convergence of our method. Extensive experiments are conducted on image and video generation to validate the superiority of the RLR. Furthermore, we propose a novel prompt technique that is natural for the RLR to achieve a synergistic effect.
Figure 1: Image and video samples generated from the model fine-tuned by our RLR optimizer. Please refer to Appendices D, E, and F for more qualitative examples.
@@ 1 Introduction
Probabilistic diffusion model (DM) (Sohl-Dickstein et al., 2015; Ho et al., 2020; Zhang et al., 2024b; Gao & Li, 2025) has emerged as a transformative framework in high-fidelity data generation, demonstrating unmatched capabilities in diverse applications such as image synthesis (Podell et al., 2023), video generation (Wang et al., 2023a), and multi-modal data modeling. These models operate by recursively denoising latent representations, as shown in Figure 3, effectively capturing complex data distributions. However, fine-tuning DMs in the post-training phase remains a daunting challenge, since the gradient estimation through the recursive structure imposes excessive computation overhead (Clark et al., 2023). This challenge has limited the broader deployment of DM in dynamic and resource-constrained environments.
Figure 2: The recursive structure of diffusion models for gradient estimation.
@@
It is a natural way to fine-tune DMs via full backpropagation (BP) through all time steps (Rumelhart et al., 1986), which is theoretically functional, providing precise gradient estimation over the entire diffusion chain. However, the computational and memory overhead of BP scales prohibitively with the model size and the number of diffusion steps (Prabhudesai et al., 2023; Yuan et al., 2024a; Clark et al., 2023; Prabhudesai et al., 2024), making full BP impractical for most real-world scenarios. Specifically, training Stable Diffusion 1.4 by full BP with a batch size of 1 and 50 time steps would require approximately 1TB of GPU RAM (Prabhudesai et al., 2023). Thus, truncating recursive differentiation becomes a common practice to alleviate memory overhead. But truncated BP suffers from structural bias, as it terminates gradient computation before sourcing to the input, only considering a limited subset of the diffusion chain. Fine-tuning based on a biased gradient estimation can inadvertently impair the optimization performance, resulting in model collapse, i.e., contents generated reducing to pure noise. Empirical evidence, in Figure 3, shows that truncated gradients result in a significant drop in reward scores during training: the fewer the truncated time steps, the more severe the model collapse. Moreover, truncated BP fails to capture the multi-scale information across all time steps due to the absence of differentiation on early steps. The hierarchical feature of generation imposes a thorough gradient evaluation to retain fidelity from pixel to structural level.
@@
Reinforcement learning (RL) (Schulman et al., 2017) as a gradient computation trick has enabled another branch of DM fine-tuning (Lee et al., 2023; Black et al., 2023; Wallace et al., 2024; Fan et al., 2024). It typically ignores the differentiable connection between steps and recovers the gradient by estimation. RL avoids caching intermediate activations, significantly reducing memory requirements. It also supports gradient computation in a divided manner under extreme circumstances to accommodate insufficient memory. The cost to pay is the high variance of the estimated gradient. Even if the estimator is unbiased, the variance can result in wild sample-inefficient updates, demonstrated by the slow convergence during training.
These limitations of BP and RL underscore the necessity of a more efficient, scalable, and stable fine-tuning approach that harmonizes computational tractability with optimization efficacy. To address this, we first investigate the recursive architecture of the DM and propose the problem of finding the minimal variance gradient estimator. Informed by perturbation-based estimation using Likelihood Ratio (LR) techniques (Jiang et al., 2024; Ren et al., 2025) (detailed reviews of LR techniques are provided in the Appendix A), we propose the Recursive Likelihood Ratio (RLR) optimizer. By utilizing the inherent noise in the DM, a local computational graph is enabled for pathwise gradient estimation as shown in Figure 4, which can reduce the variance and better capture the multi-scale information. The RLR estimator is unbiased and has lower variance under the same computation budgets as other methods. Through perturbation-based computational graph rearrangement, the RLR mitigates the structural bias of truncated BP and the high variance of RL, capable of better capturing multi-scale visual information. Our optimizer shares similarity with zeroth-order optimizer since both need perturbation to estimation gradient, but the RLR utilize a local BP chain to enable low-variance estimation. Therefore, we name it as Half-Order optimizer. Our contributions are threefold:
  • We provide a systematic analysis of gradient estimation in DMs, identifying a structured design space of estimators. Then, we formulate the RLR optimizer in the design space under the protocol of minimizing variance with a limited computational budget.
@@
  • Extensive evaluation of Text2Image and Text2Video tasks are conducted. RLR consistently achieves higher reward scores across multiple human preference reward models and outperforms SOTA video models on the VBench benchmark. Furthermore, we propose a prompt technique for our RLR, validating the intuition and applicability of our method.
  • We conduct a rigorous theoretical analysis of RLR’s design, proving its unbiasedness, bounding its estimator variance, and establishing convergence guarantees. These results formally justify the empirical success of RLR and explain the deficiency of prior methods.
@@
2 Related Work
Diffusion probabilistic models have achieved state-of-the-art performance in multi-modal generation (Ho et al., 2020; Rombach et al., 2022). However, aligning pre-trained DMs for downstream tasks remains challenging. RL and supervised post-training have been widely explored to incorporate human preferences or task-specific objectives, with recent methods including RLHF-inspired approaches, DPO variants, and reward-model-guided fine-tuning (Ziegler et al., 2019; Fan et al., 2024; Wallace et al., 2024; Prabhudesai et al., 2023). Yet, RL-based fine-tuning often suffers from high variance and sample inefficiency, while truncated BP reduces memory cost but introduces structural bias that can lead to model collapse(Prabhudesai et al., 2023; 2024; Xu et al., 2024b). Recent forward-learning methods based on stochastic gradient estimation offer a promising alternative with lower computational and memory costs (Salimans et al., 2017; Peng et al., 2022; Chen et al., 2023). Please refer to Appendix A for extended related works.
3 Unbiased Minimal Variance Gradient Estimator for Diffusion Model
3.1 Problem Formulation
DMs generate data by transforming a noise sample $x_{T}$ into a data sample $x_{0}$ through a recursive denoising process, as shown in Figure 3. This generation proceeds through $T$ steps of stochastic updates, each governed by a parameterized mapping $\phi_{t}$, where the implementation of $\phi_{t}$ may vary across steps, yielding the transformation:
$x_{t-1}=\phi_{t}(x_{t},z_{t};\theta)$ (1)
where $\theta$ denotes the model parameter, and $z_{t}=\sigma_{t}\epsilon_{t},\epsilon_{t}\sim\mathcal{N}(\bm{0},I)$ is a stochastic perturbation. The noise term $z_{t}$ can be either the inherent noise in the DM, i.e., $x_{t-1}=\varphi(x_{t};\theta)+z_{t}$, or the injected noise to the parameter, i.e., $x_{t-1}=\varphi(x_{t};\theta+z_{t})$. The function $\varphi$ denotes the shared backbone network of the DM, reused across different steps, while $\phi_{t}$ should be used with the subscript $t$ to indicate which time-step it operates on. The full generation chain can be represented recursively as:
$x_{0}=\phi_{1:T}(x_{T},z_{1:T};\theta)=\phi_{1}\circ\phi_{2}\circ\dots\circ\phi_{T}(x_{T},z_{1:T};\theta)=$ (2)
$\phi_{1}(\phi_{2:T}(x_{T},z_{2:T};\theta),z_{1};\theta)=\phi_{1}(\phi_{2:T-1}(\phi_{T}(x_{T},z_{T};\theta),z_{2:T-1};\theta),z_{1};\theta),$
where $z_{1:T}$ are independent noise terms injected at each step. Once a pre-trained model is available, a reward model $R(x_{0})$ is introduced to guide post-training towards generating samples of higher semantic or perceptual quality. The resulting fine-tuning objective becomes:
$\max_{\theta}\mathbb{E}[R(x_{0})]=\max_{\theta}\mathbb{E}_{z_{1:T}}[R(\phi_{1:T}(x_{T},z_{1:T};\theta))].$ (3)
Fine-tuning DMs through the recursive structure poses a critical challenge: how to efficiently and accurately estimate the gradient of the expected reward with respect to model parameters. A natural objective is to construct a gradient estimator for $\mathbb{E}[R(x_{0})]$ that is unbiased and has minimal variance, under a given computational and memory budget. We formulate the problem as the following constrained optimization problem over the space of gradient estimators:
$\min_{G\in\mathcal{G}}\quad\operatorname{Var}(G)\qquad\text{s.t.}\quad\nabla_{\theta}\mathbb{E}[R(x_{0})]=\mathbb{E}[G],\quad\mathcal{C}(G)\leq\mathcal{B},$ (4)
where $\mathcal{G}$ denotes the space of all unbiased gradient estimators $G$ for the objective $\mathbb{E}[R(x_{0})]$, with the sample $x_{0}$ being obtained via a generative chain defined by $x_{0}=\phi_{1:T}(x_{T};z_{1:T},\theta)$. The term $\operatorname{Var}(G)$ refers to the variance of the estimator, which we aim to minimize. The cost function $\mathcal{C}(G)$ quantifies the total computational and memory overhead incurred by the estimator $G$, typically measured in terms of the length of backward passes or the volume of intermediate activations stored. Finally, $\mathcal{B}$ represents the computational budget available for gradient estimation.
3.2 $\mathcal{G}$: The Unbiased Gradient Estimator Design Space in Diffusion Models
We now characterize the feasible design space $\mathcal{G}$ for unbiased gradient estimators in DMs. Due to the recursive structure of the generative chain, the gradient estimator must propagate through all $T$ steps. At each step $t$, we may choose one of three gradient estimation strategies:
  • First-Order (FO) uses exact backpropagation through $\phi_{t}$. Zeroth-Order (ZO) perturbs the parameters directly, e.g., $\varphi_{t}(x_{t};\theta+\sigma_{t}\epsilon_{t})$, and estimates the gradient via $\frac{R(\varphi(\cdot;\theta+\sigma_{t}\epsilon_{t}))}{\sigma_{t}}\epsilon_{t}$, based solely on function evaluations (Salimans et al., 2017).
  • We propose unbiased Half-Order (HO) gradient estimator which utilizes the inherent noise $z_{t}$ (rather than extrinsic perturbation) and applies the Likelihood Ratio technique, producing an estimator of the form: $R(x_{0})\cdot D_{\theta}^{\top}\phi_{t:t+h-1}\cdot\nabla\log f(z_{t})$, where $D_{\theta}\phi_{t:t+h-1}$ denotes the Jacobian of a local sub-chain of length $h$, and $f(\cdot)$ denotes the noise density. The HO allows a $h$-length sub-chain, starting from any $t$ in the diffusive chain. RL is a special case of HO method with $h=1$.
The full estimator design space is thus defined as:
$\mathcal{G}_{\text{full}}:=\{G=(g_{1},\dots,g_{T})\mid g_{t}\in\{\text{FO},\text{HO},\text{ZO}\}\;\;\forall 1\leq t\leq T\},$ (5)
Table 1: Comparison of gradient estimators.
Method Unbiasedness Variance Memory